OpenStudy (anonymous):

i just need help with the formula here...find the number of ways in which each committee can be selected. 3 people from a group of 5

OpenStudy (anonymous):

5C3

OpenStudy (lgbasallote):

the formula would be combinations \[nCr = \frac{n!}{r!(n-r)!}\] does that help?

OpenStudy (anonymous):

so well thats what i did...5!/(2!-3!)...and the answer was wrong

OpenStudy (lgbasallote):

that's because it's \[\frac{5!}{3!2!}\]not minus

OpenStudy (amistre64):

5P3 ---- 3!

OpenStudy (amistre64):

5.4.3 5.4 ----- = --- = 20/2 3.2.1 2

OpenStudy (anonymous):

oo okay oops!...thanks:)